Sol Summary Report – February 14th



Crew 275 Sol Summary Report 14Feb2023

Sol: 2

Summary Title: Journey to the poles.

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary: Today the crew was surprised by the snow. With the solar panels covered by a blanket of a few inches/centimeters, we had to pay even more attention than usual to our power consumption. Before real-life deployment during a dedicated EVA, the crew trained to build a first device dedicated to atmospheric measurements in the RAM. However, the temperature became too cold for long activities to be conducted in there (2°C/36°F). We limited navigation between the different modules as much as possible, as snow started to also cover the floor in the different tunnels. As a result, most of the crew had to adapt their activities and gathered in the Hab to work from there, which made the atmosphere a bit more noisy than usual.

However, we remain productive and managed to calibrate the first sensors that will monitor the environmental conditions within the station. They are now connected to the local network. In the GreenHab, the activities were successfully conducted, and we now have a fully operational aquaponics system, containing 28 plants and 8 fishes. As for the astronomy project, even though the robotic observatory has been fixed yesterday, no observations were performed during the night (both at the MDRS and at the other observatory in New Mexico).

Look Ahead Plan:

The weather conditions have already impacted the schedule of the different activities and it is to be expected that this impact will extend to the next few days. If the land remains covered with snow, it may not be possible to proceed with the deployment of at least one of the two devices dedicated to atmospheric studies (LOAC). Depending on the width of snow at the targeted observation site on Thursday, the whole EVA may have to be postponed for safety reasons.

Some temperature sensors have been installed in the Science Dome and on the Lower Deck of the Hab. They are connected to the local network and will record data every few minutes during the whole night so that tomorrow we wan check if there were any problems, in particular regarding network.

Anomalies in work: Nothing to report.

Weather: Snowy, low -5°C/23°F, high 2°C/35°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, Operations Report, Journalist Report, HSO Report, GreenHab Report, Astronomy Report.

Support Requested:

Requesting the possibility to add the following email address to the Mission Support mailing list (we apologize for not already including in the previous request):


Copyright © The Mars Society. All rights reserved. | Main Site